Interpret the utility function: U(x1 , x2,...,xn )= β1x1 + β2x2 + ··· + βnxn ; Similarly, One study reports a utility function that had the form : U(TW,TT,C)= −0.147TW − 0.0411TT − 2.24C, where TW = total walking time to and from bus or car; TT = total time of trip in minutes, C = total cost of trip in dollars; rigorously interpret the meaning of that utility function. 


1
Expert's answer
2021-01-11T12:00:52-0500

Utility function "U(x1,x2,...,xn)" shows the measure of satisfaction that consumers receive for choosing of some set of goods. The utility function "U(TW,TT,C)=-0.147TW-0.0411TT-2.24C" is always less than zero (an individual does not like any trips). This function shows the pindividual utility as a preferences over a set of the following actions during a trip: trip from some start point to some finish point. Individual can chose within 2 options: walking (denoted by TW) or using a bus/car during a period that is equal to "TT-TW" (incliding payment for the traveling costs denoted by C).

The first order derivatives of the U with respect to TW, TT, and C can relate a change in U to a change in arguments TW, TT, and C).

"dU\/dTW=-0.147<0" (1)

"dU\/dTT=-0.0411<0" (2)

"dU\/dC=-2.24<0" (3)

The inequations (1) and (2) shows that individual does not like to go from a home to bus/car rather than to use car/bus. Hovewer. individual does not like to pay.

Let's assume that the TW is equal to zero and analyze the indifference curve based on the U1 and U2: "U1(TT1,C1)" = "U2(TT2,C2)."

Hence, "-0.0411TT1-2.24C1=-0.0411TT2-2.24C2"

Then, "-0.0411(TT1-TT2)=-2.24(C2-C1)"

Hence, if the total time of trip increases, then cost of trip should be decresed to provide the constant utility.

Conclusions:

1) An individual does not like any trips

2) An individual does not like to go from a home to bus/car rather than to use car/bus

3) Positive total walking time to and from bus or car is result of inequation (3) that shows that utility decreases when the cost for trip increases.